A ble with a marble r , linear,speed tv. energy-is o What al KE uPpl cst 0 0.4m Vertically-Resa ーis.FAA.ーなnear-speed asZA.balLa protees ra man A 60 フ3.5pm
Answer #1

1.

Moment of inertia of marble

I = (2/5) mr^2 = 0.4 mr^2

Rotational KE

KEr = 0.5 Iw^2 = 0.5 ( 0.4 mr^2) ( v/r) ^2

KEr = 0.2 mv^2 (I)

Total kinetic energy

KET = 0.5 mv^2 + 0.5 I w^2 = 0.5 mv^2 + 0.5 (0.4 mv^2)

KET = 0.7 mv^2 (ii)

Considering equ (I) and (ii)

KEr/ KET = 0.2/0.7

KEr = (2/7) KET

============

Plz post other problems seperately,

Do comment in case any doubt, will reply for sure, Goodluck